For two events, A and B, P(A=0.2, P(B)=0.50, and P(A|B)=0.2.

a. Find P(A∩B) .b. Find P(B|A). (Simplify your answer)

b. P(B|A)=__________(Simplify your answer.)
Answer #1

Given that, P(A) = 0.2, P(B) = 0.50 and P(A | B) = 0.2

we want to find, the following probabilities,

a) We know,

P(B)

P(A cap B) = P(A|B) * P(B)

P(A n B) = 0.2 * 0.50

> P(An B) = 0.1

b)

P (BA) = P(A)

0.1

P(B|A) = 0.5